# On elliptic and quasiregularly elliptic manifolds

### v1

@@ 1.5. Are closed quasiregularly elliptic manifolds elliptic?
This would be the case if every such manifold were BLD-elliptic, since every BLD map is both quasiregular and Lipschitz. There is not much evidence for this conjecture; Gromov [13, 6.42] sketches a proof, further discussed by Bonk and Heinonen [7, §2], that if there is a quasiregular map $\mathbb{R}^{n}\to M$, then in fact there is a Hölder quasiregular map [^5] The case $M=S^{n}$ was done by Miniowitz [26]., but it’s not clear how one would upgrade this to a Lipschitz quasiregular map.
In general, a quasiregular map need not be even locally Lipschitz: for example, consider the radial stretch map $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined by $f(x)=|x|^{\alpha}\cdot x$, for $\alpha<1$. Moreover, there are examples of quasiregular maps with wild branch sets which are not even topologically conjugate to Lipschitz maps [16]. On the other hand, if one assumes that the branch set (the subset of the domain at which the map is not a local homeomorphism) is contained in a subcomplex of codimension $2$, then the map is always topologically conjugate to a locally Lipschitz one, namely a PL branched cover, due to work of Luisto and Prywes [22]. Thus, even if one wants to show that quasiregularly elliptic manifolds are BLD-elliptic—a more restrictive condition than Gromov ellipticity—the main obstacle is still likely to be the global, rather than local, geometry of the map.
@@ 1.6. Are closed elliptic manifolds quasiregularly elliptic?
Gromov suggested in [13, p. 67] that it might be possible to obtain a quasiregular map by somehow optimizing over all Lipschitz maps of positive asymptotic degree: for example, looking at $1$-Lipschitz maps which attain the highest possible degree on every $R$-ball. The example of $S^{2}$ is illustrative. Here, for any choice of rectangle on the equator one gets a quasiregular map which is a cover branched at the preimages of the vertices of the rectangle and sends alternate squares of a checkerboard to the north and south hemisphere. If one gives the squares side length $\pi$, then this map can be made $1$-Lipschitz, and presumably has the highest possible degree on all balls. On the other hand, one can achieve the same degree on balls if one forces two branching points to collide with each other, at which point entire lines of the checkerboard are mapped to the point of collision, and the map is no longer quasiregular. Moreover, at that point one can start moving whole strips of the checkerboard laterally, giving many more degrees of freedom of travel through maps that do not seem to even be close to a quasiregular map. In other words, we have described a large continuous family of optimal maps, of which a non-dense subset is quasiregular.
@@ 1.7. Key examples for future work
    Let $\Sigma$ be an $n$-dimensional simply connected rational homology sphere. For what $12\leq k\leq n-$ 1 is $M=(T^{n-k}\times S^{k})\mathbin{\#}\Sigma$ elliptic or quasiregularly elliptic? This series of examples interpolates between those of Theorem 1.4 and the rational invariance of Theorem 1.5. Note that $M$ cannot be a nilpotent space, as the connected sum gives a copy of the module $H_{p}(\Sigma)[\mathbb{Z}^{n-k}]$ in $\pi_{p}(M)$, where $p$ is the minimal degree such that $H_{p}(\Sigma)\neq 0$.
@@
    Consider the simplest example of a simply connected non-formal manifold mentioned above. This is the $7$-manifold $M$ which completes the pullback square
@@
    Is this manifold elliptic? We suspect that it is not, but do not know how to prove this.
@@
    Nevertheless, it seems difficult and perhaps impossible to get rid of the logarithmic factor.
@@
    One such space is constructed in [5, Remark 5.2] as follows. Take the real Poincaré duality space
@@
    where * is a basepoint. The cup product $H^{2}(Y)\times H^{2}(Y)\to H^{4}(Y)$ is the quadratic form $\langle 2,1,1,-1,-1,-1\rangle$, which has discriminant $-2$, and therefore is not rationally equivalent to the quadratic form induced by the cup product $H^{2}(T^{4})\times H^{2}(T^{4})\to H^{4}(T^{4})$. Embed $Y$ in $\mathbb{R}^{10}$. The boundary of a thickening of this embedding is a $9$-manifold with the desired property.
@@ 2. Proof of the cohomological obstruction
In this section we prove Theorem 1.2. We start with the elliptic version, first given in [5].
@@
For the elliptic case, we use flat differential forms as introduced by Whitney [37, Ch. IX]. These can be thought of as formal limits of smooth forms in the flat norm given by
@@
Flat forms have a number of relevant properties, for example the pullback of a smooth (or flat) form along a Lipschitz function is a flat form [10, Theorem 3.6]. Following Whitney, we write $\Omega_{\flat}^{*}(M)$ for the complex of flat forms on $M$.
@@
Since flat limits preserve wedge products [37, §IX.17] and differentials (by definition), this implies that the ring homomorphism $\mathbb{R}[w_{1},\ldots,w_{r}]\to\Omega^{*}_{\flat}(\mathbb{R}^{n})$ defined by $w_{i}\mapsto\omega_{i,\infty}$ passes to a well-defined map on the quotient ring $H^{*}(M,\mathbb{R})\cong\mathbb{R}[w_{1},\ldots,w_{r}]/(p_{1},\ldots,p_{s})$, giving a ring homomorphism
@@
The proof of the quasiregular case uses a similar overall strategy, but is somewhat more complicated, using ideas from [31]. The main idea is still to compare forms defined on larger and larger balls, so that in the limit primitives go to zero. However, the scaling factor is no longer related to the size of the ball; instead, the existence of a useful scaling factor is guaranteed by quasiregularity. Moreover, the sequence converges in an $L^{p}$ rather than a flat sense.
@@
be the counting function for $f$. We consider the pullbacks by $f$ of the differential forms $\omega_{i}$ and $\alpha_{j}$. The pullbacks have to be rescaled, following the procedure in [31, Section 4]. If $B_{k}\subset\mathbb{R}^{n}$ is a sequence of balls, we define forms on $B_{1}=B(0,1)$ by
@@
where $p_{i}=n/a_{i}$ and $q_{j}=n/b_{j}$ and $T_{k}(x)=x_{k}+r_{k}x$ when $B_{k}=B(x_{k},r_{k})$. We summarize the main results in [31, Section 4] as applied to our setting.
@@
    The sequence $\alpha_{j,k}$ converges strongly in $L^{q_{i}}(B_{1})$ to $\alpha_{j,\infty}$ and $\alpha_{j,\infty}\in W^{1,q_{i}}(B_{1})$.
@@
If $a_{i}=n$, then $\omega_{i}=d\operatorname{vol}_{M}$ and $\omega_{i,\infty}=d\operatorname{vol}_{\mathbb{R}^{n}}$ on $B_{1}$.
@@
We first show that $\alpha_{j,k}$ converges to 00:
@@
which converges to 00 as $k\to\infty$ by Lemma 2.2.
@@
[…]
using integration by parts. This term converges to 00 as $k\to\infty$ by the strong convergence of $\alpha_{j,k}$.
@@
Therefore the sequence is bounded and when $b_{j}\neq n$ there exists a subsequence that converges weakly. Since the sequence converges to 00 in distribution we arrive at the claim of the lemma. ∎
@@
Suppose $\omega_{k}$ and $\omega_{k}^{\prime}$ are two sequences of $a$-forms and $a^{\prime}$-forms constructed as above, where $a$ and $a^{\prime}$ do not equal 00. Let $p=n/a$ and $p^{\prime}=n/a^{\prime}$. If $\omega_{k}\to\omega_{\infty}$ weakly and $\omega_{k}^{\prime}\to\omega_{\infty}^{\prime}$ weakly in $L^{p}(B_{1})$ and $L^{p^{\prime}}(B_{1})$, repectively, then $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{k}\wedge\omega_{k}^{\prime}$ weakly in $L^{n/(a+a^{\prime})}(B_{1})$, if $a+a^{\prime}<n$, and converges in distribution, if $a+a^{\prime}=n$.
@@
where $q^{-1}=1-p^{-1}-(p^{\prime})^{-1}$. This converges to 00 by the strong convergence of $\tau_{k}$ and the boundedness of $\omega_{k}^{\prime}$ in $L^{p^{\prime}}$, where both properties are shown in Lemma 2.2. For the second term, $d\phi\wedge\tau_{\infty}$ is an element of $L^{p^{\prime}/(p^{\prime}-1)}$ since $a+b\leq n$. By the weak convergence of $\omega_{k}^{\prime}$, this term also converges to 00. This gives that $\omega_{k}\wedge\omega_{k}^{\prime}$ converges to $\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution. If $a+b<n$, by Hölder’s inequality we know that
@@
We next must show that $\varphi_{\infty}$ is injective. Let $\omega$ and $\omega^{\prime}$ be representatives of Poincaré dual classes on $M$. By Lemma 2.4, $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution, where the limit is in $L^{1}(B_{1})$. Let $V$ be a positive representative of the fundamental class of $M$. Then $\omega\wedge\omega^{\prime}-V=d\alpha$ and, by Lemma 2.3, $d\alpha_{k}\to 0$ in distribution. It therefore suffices to show that $V_{k}=\frac{1}{A(B_{k})}T_{k}^{*}f^{*}V$ does not converge to 00.
@@
by Lemma 2.2, part (iv). Thus, $V_{k}$ does not converge to 00 in distribution, which proves the claim.
@@ 3. Consequences of the cohomological obstruction
If $M$ is a quasiregularly elliptic (or Gromov-elliptic) closed $n$-manifold, then $\pi_{1}(M)$ is virtually abelian.
@@
By Theorem 1.1, $\pi_{1}(M)$ is virtually nilpotent. Since covers of (quasiregularly, Gromov/̄)elliptic spaces are elliptic, we can assume by passing to a finite cover that $\pi_{1}(M)$ is nilpotent and torsion-free [8, Lemma 2.42].
Let $\Gamma=\pi_{1}(M)$, and let $G$ be its Mal’cev completion: a simply connected (and therefore contractible) nilpotent Lie group which contains it as a lattice [23]. Then the nilmanifold $N=G/\Gamma$ is a $K(\Gamma,1)$, so there is a map $g:M\to N$ which induces a $\pi_{1}$-isomorphism. The pullback and Theorem 1.2 induce maps
[…]
whose composition is injective on $H^{1}$, and therefore on the subalgebra generated by $H^{1}$. Therefore the subalgebra of $H^{*}(N;\mathbb{R})$ generated by $H^{1}(N;\mathbb{R})$ is isomorphic to $\Lambda^{*}\mathbb{R}^{k}$ where $k=\operatorname{rk}H^{1}(N)$. We will show that this implies that $\Gamma$ is abelian.
Nomizu’s theorem [27] implies that $H^{*}(N;\mathbb{R})$ is isomorphic as a ring to the cohomology of the differential graded algebra (DGA) $\Omega^{*}(G)^{G}$ of $G$-invariant differential forms on $G$. We denote this cohomology by $H^{*}(G)$.
@@
Now consider the pullback of $H^{*}(T^{k})$ along $f$. Since $H^{*}(\tilde{M};\mathbb{R})$ embeds in $\Lambda^{*}\mathbb{R}^{n}$, the image of this pullback is the subalgebra generated by a $k$-dimensional vector subspace of $\Lambda^{1}\mathbb{R}^{n}$. This subalgebra is isomorphic to $\Lambda^{*}\mathbb{R}^{k}$, meaning that $f^{*}$ is injective. Dually, $f$ is surjective on homology. ∎
@@
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold. Then $\pi_{1}(M)$ cannot have rank exactly $n-1$.
@@
Suppose $\pi_{1}(M)$ has rank $n-1$. By passing to a finite cover, we can assume that $\pi_{1}(M)\cong\mathbb{Z}^{n-1}$. Consider an injective map $i:H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}$. The image of $i$ contains an $(n-1)$-dimensional subspace of $\Lambda^{1}\mathbb{R}^{n}$, without loss of generality $\langle dx_{1},\ldots,d_{n-1}\rangle$. The product $dx_{1}\wedge\cdots\wedge dx_{n-1}$ must pair nontrivially with a $1$-form, which then has to contain a nonzero $dx_{n}$ term. This is a contradiction. ∎
@@
Suppose that $A\subset\Lambda^{*}\mathbb{R}^{n}$ is a graded subalgebra which satisfies Poincaré duality, and let $x\in A^{1}$. Then $A/(x)$ also satisfies Poincaré duality.
@@
Otherwise, $a=x_{1}a^{\prime}$ for some $a^{\prime}\in\Lambda^{*}\mathbb{R}^{n}$ which is not necessarily in $A$. We show that this cannot happen. Consider the vector space $V$ of elements of $A^{k}$ that have this form. We can choose a vector space $W\subset A^{n-k}$ such that the wedge product defines a nondegenerate bilinear form on $V\otimes W$. In particular, multiplication by $x_{1}$ defines a bijection $W\to x_{1}W$. Now we can similarly build a vector space $V^{\prime}\subset A^{k-1}$ which is Poincaré dual to $x_{1}W$. The map $V^{\prime}\to V$ defined by multiplication by $x_{1}$ is injective, and hence surjective since $V^{\prime}$ and $V$ have the same dimension. Therefore $a$ is contained in the ideal $(x)$ and $\bar{a}=0$, contradicting our assumption. ∎
@@
Let $M$ be such a manifold. By passing to a finite cover we may assume that $M$ has fundamental group $\mathbb{Z}^{2}$ or $\mathbb{Z}$, or is $T^{4}$ by Theorem 1.4. If it is not $T^{4}$, then from Proposition 3.4 it follows that $M$ has the same real cohomology ring as $T^{2}\times S^{2}$ or $S^{1}\times S^{3}$, and in particular has Euler characteristic zero. By [17, Corollary 6.11.1], $M$ is determined by the total Stiefel–Whitney class of the bundle $S^{2}\to M\to T^{2}$ or $S^{3}\to M\to S^{1}$, respectively. Since this class is of finite order, $M$ is finitely covered by either $T^{2}\times S^{2}$ or $S^{1}\times S^{3}$. ∎
This completes the classification of elliptic and quasiregularly elliptic $4$-manifolds up to finite covers. If $\pi_{1}(M)$ is finite, then $M$ is finitely covered by a simply connected manifold which must come from a finite list of 19 manifolds given in [15, Appendix A].

### v3

@@ 1.5. Are closed quasiregularly elliptic manifolds elliptic?
This would be the case if every such manifold were BLD-elliptic, since every BLD map is both quasiregular and Lipschitz. There is not much evidence for this conjecture; Gromov [GrMS, 6.42] sketches a proof, further discussed by Bonk and Heinonen [BH, §2], that if there is a quasiregular map $\mathbb{R}^{n}\to M$, then in fact there is a Hölder quasiregular map [^4] The case $M=S^{n}$ was done by Miniowitz [Mini]., but it’s not clear how one would upgrade this to a Lipschitz quasiregular map.
In general, a quasiregular map need not be even locally Lipschitz: for example, consider the radial stretch map $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined by $f(x)=|x|^{\alpha}\cdot x$, for $\alpha<1$. Moreover, there are examples of quasiregular maps with wild branch sets which are not even topologically conjugate to Lipschitz maps [HR]. On the other hand, if one assumes that the branch set (the subset of the domain at which the map is not a local homeomorphism) is contained in a subcomplex of codimension $2$, then the map is always topologically conjugate to a locally Lipschitz one, namely a PL branched cover, due to work of Luisto and Prywes [LuPr]. Thus, even if one wants to show that quasiregularly elliptic manifolds are BLD-elliptic—a more restrictive condition than Gromov ellipticity—the main obstacle is still likely to be the global, rather than local, geometry of the map.
@@ 1.6. Are closed elliptic manifolds quasiregularly elliptic?
Gromov suggested in [GrMS, p. 67] that it might be possible to obtain a quasiregular map by somehow optimizing over all Lipschitz maps of positive asymptotic degree: for example, looking at $1$-Lipschitz maps which attain the highest possible degree on every $R$-ball. The example of $S^{2}$ is illustrative. Here, for any choice of rectangle on the equator one gets a quasiregular map which is a cover branched at the preimages of the vertices of the rectangle and sends alternate squares of a checkerboard to the north and south hemisphere. If one gives the squares side length $\pi$, then this map can be made $1$-Lipschitz, and presumably has the highest possible degree on all balls. On the other hand, one can achieve the same degree on balls if one forces two branching points to collide with each other, at which point entire lines of the checkerboard are mapped to the point of collision, and the map is no longer quasiregular. Moreover, at that point one can start moving whole strips of the checkerboard laterally, giving many more degrees of freedom of travel through maps that do not seem to even be close to a quasiregular map. In other words, we have described a large continuous family of optimal maps, of which a non-dense subset is quasiregular.
@@ 1.7. Key examples for future work
    Are the examples of non-nilpotent elliptic manifolds given in Theorem 1.6 also quasiregularly elliptic?
@@
    The simplest example of a simply connected non-formal manifold is the $7$-manifold $M$ which completes the pullback square
@@
    Is this manifold elliptic? We suspect not, but do not know how to prove this.
@@
    But the logarithmic factor seems difficult, and perhaps impossible, to get rid of.
@@
    One such space is constructed in [BeGM, Remark 5.2] as follows. Take the real Poincaré duality space
@@
    where * is a basepoint. The cup product $H^{2}(Y)\times H^{2}(Y)\to H^{4}(Y)$ is the quadratic form $\langle 2,1,1,-1,-1,-1\rangle$, which has discriminant 2-2, and therefore is not rationally equivalent to the quadratic form induced by the cup product $H^{2}(T^{4})\times H^{2}(T^{4})\to H^{4}(T^{4})$. Embed $Y$ in $\mathbb{R}^{10}$. The boundary of a thickening of this embedding is a $9$-manifold with the desired property.
@@ 2. Proof of the cohomological obstruction
In this section we prove Theorem 1.2. We start with the elliptic version, first given in [BeGM].
@@
For the elliptic case, we use flat differential forms as introduced by Whitney [Whi, Ch. IX]. These can be thought of as formal limits of smooth forms in the flat norm given by
@@
Flat forms have a number of relevant properties, for example the pullback of a smooth (or flat) form along a Lipschitz function is a flat form [GKS, Theorem 3.6]. Following Whitney, we write $\Omega_{\flat}^{*}(M)$ for the complex of flat forms on $M$.
@@
Since flat limits preserve wedge products [Whi, §IX.17] and differentials (by definition), this implies that the ring homomorphism $\mathbb{R}[w_{1},\ldots,w_{r}]\to\Omega^{*}_{\flat}(\mathbb{R}^{n})$ defined by $w_{i}\mapsto\omega_{i,\infty}$ passes to a well-defined map on the quotient ring $H^{*}(M,\mathbb{R})\cong\mathbb{R}[w_{1},\ldots,w_{r}]/(p_{1},\ldots,p_{s})$, giving a ring homomorphism
@@
The proof of the quasiregular case uses a similar overall strategy, but is somewhat more complicated, using ideas from [Pry]. The main idea is still to compare forms defined on larger and larger balls, so that in the limit primitives go to zero. However, the scaling factor is no longer related to the size of the ball; instead, the existence of a useful scaling factor is guaranteed by quasiregularity. Moreover, the sequence converges in an $L^{p}$ rather than a flat sense.
@@
be the counting function for $f$. We consider the pullbacks by $f$ of the differential forms $\omega_{i}$ and $\alpha_{j}$. The pullbacks have to be rescaled, following the procedure in [Pry, Section 3]. If $B_{k}\subset\mathbb{R}^{n}$ is a sequence of balls, we define forms on $B_{1}=B(0,1)$ by
@@
where $p_{i}=n/a_{i}$ and $q_{j}=n/b_{j}$ and $T_{k}(x)=x_{k}+r_{k}x$ when $B_{k}=B(x_{k},r_{k})$. We summarize the main results in [Pry, Section 3] as applied to our setting.
@@
    The sequence $\alpha_{j,k}$ converges strongly in $L^{q_{i}}(B_{1})$ to $\alpha_{j,\infty}\in W^{1,q_{i}}(B_{1})$.
@@
If $a_{i}=n$, then $\omega_{i}=d\operatorname{vol}_{M}$ and $\int_{B_{1}}\omega_{i,\infty}=\operatorname{vol}(B_{1})$.
@@
We first show that $\alpha_{j,k}$ converges to 0:
@@
which converges to 0 as $k\to\infty$ by Lemma 2.2.
@@
$\left|{\int_{B_{1}}d\alpha_{j,k}\wedge\phi}\right|=\left|{-\int_{B_{1}}\alpha_{j,k}\wedge d\phi}\right|,$
using integration by parts. This term converges to 0 as $k\to\infty$ by the strong convergence of $\alpha_{j,k}$.
@@
Therefore the sequence is bounded and when $b_{j}\neq n$ there exists a subsequence that converges weakly. Since the sequence converges to 0 in distribution we arrive at the claim of the lemma. ∎
@@
Suppose $\omega_{k}$ and $\omega_{k}^{\prime}$ are two sequences of $a$-forms and $a^{\prime}$-forms constructed as above, where $a$ and $a^{\prime}$ do not equal 0. Let $p=n/a$ and $p^{\prime}=n/a^{\prime}$. If $\omega_{k}\to\omega_{\infty}$ weakly and $\omega_{k}^{\prime}\to\omega_{\infty}^{\prime}$ weakly in $L^{p}(B_{1})$ and $L^{p^{\prime}}(B_{1})$, repectively, then $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{k}\wedge\omega_{k}^{\prime}$ weakly in $L^{n/(a+a^{\prime})}(B_{1})$, if $a+a^{\prime}<n$, and converges in distribution, if $a+a^{\prime}=n$.
@@
where $q^{-1}=1-p^{-1}-(p^{\prime})^{-1}$. This converges to 0 by the strong convergence of $\tau_{k}$ and the boundedness of $\omega_{k}^{\prime}$ in $L^{p^{\prime}}$, where both properties are shown in Lemma 2.2. For the second term, $d\phi\wedge\tau_{\infty}$ is an element of $L^{p^{\prime}/(p^{\prime}-1)}$ since $a+b\leq n$. By the weak convergence of $\omega_{k}^{\prime}$, this term also converges to 0. This gives that $\omega_{k}\wedge\omega_{k}^{\prime}$ converges to $\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution. If $a+b<n$, by Hölder’s inequality we know that
@@
We next must show that $\varphi_{\infty}$ is injective. Let $\omega$ and $\omega^{\prime}$ be representatives of Poincaré dual classes on $M$. By Lemma 2.4, $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution, where the limit is in $L^{1}(B_{1})$. Let $V$ be a positive representative of the fundamental class of $M$. Then $\omega\wedge\omega^{\prime}-V=d\alpha$ and, by Lemma 2.3, $d\alpha_{k}\to 0$ in distribution. It therefore suffices to show that $V_{k}=\frac{1}{A(B_{k})}T_{k}^{*}f^{*}V$ does not converge to 0.
@@
by Lemma 2.2, part (iv). Thus, $V_{k}$ does not converge to 0 in distribution, which proves the claim.
@@ 3. Consequences of the cohomological obstruction
If $M$ is a Gromov-elliptic or quasiregularly elliptic closed $n$-manifold, then $\pi_{1}(M)$ is virtually abelian.
@@
By Theorem 1.1, $\pi_{1}(M)$ is virtually nilpotent. Since covers of (quasiregularly, Gromov-)elliptic spaces are elliptic, we can assume by passing to a finite cover that $\pi_{1}(M)$ is nilpotent and torsion-free [CSDA, Lemma 2.42].
Let $\Gamma=\pi_{1}(M)$, and let $G$ be its Mal’cev completion: a simply connected (and therefore contractible) nilpotent Lie group which contains it as a lattice [Mal]. Then the nilmanifold $N=G/\Gamma$ is a $K(\Gamma,1)$, so there is a map $g:M\to N$ which induces a $\pi_{1}$-isomorphism. The pullback and Theorem 1.2 induce maps
(1) $H^{*}(N;\mathbb{R})\xrightarrow{g^{*}}H^{*}(M;\mathbb{R})\xhookrightarrow{\iota}\Lambda^{*}\mathbb{R}^{n}$
whose composition is injective on $H^{1}$, and therefore on the subalgebra generated by $H^{1}$. Therefore the subalgebra of $H^{*}(N;\mathbb{R})$ generated by $H^{1}(N;\mathbb{R})$ is isomorphic to $\Lambda^{*}\mathbb{R}^{k}$ where $k=\operatorname{rk}H^{1}(N)$. We will show that this implies that $\Gamma$ is abelian. Note that if $\iota g^{*}$ were induced by a homomorphism $\mathbb{Z}^{n}\to\Gamma$, this would follow from a theorem of Stallings [Sta].
Nomizu’s theorem [Nomizu] implies that $H^{*}(N;\mathbb{R})$ is isomorphic as a ring to the cohomology of the differential graded algebra (DGA) $\Omega^{*}(G)^{G}$ of $G$-invariant differential forms on $G$. We denote this cohomology by $H^{*}(G)$.
@@
[…]
@@
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold. Then $\pi_{1}(M)$ cannot have rank exactly n1n-1.
@@
Suppose $\pi_{1}(M)$ has rank n1n-1. By passing to a finite cover, we can assume that $\pi_{1}(M)\cong\mathbb{Z}^{n-1}$. Consider an injective map $i:H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}$. The image of $i$ contains an (n1)(n-1)-dimensional subspace of $\Lambda^{1}\mathbb{R}^{n}$, without loss of generality $\langle dx_{1},\ldots,d_{n-1}\rangle$. The product $dx_{1}\wedge\cdots\wedge dx_{n-1}$ must pair nontrivially with a $1$-form, which then has to contain a nonzero $dx_{n}$ term. This is a contradiction. ∎
@@
Suppose that $A\subset\Lambda^{*}\mathbb{R}^{n}$ is a graded subalgebra which satisfies Poincaré duality, and let $x\in A^{1}$. Then $A/(x)$ also satisfies Poincaré duality, and moreover $A/(x)$ and $(x)$ have the same dimension.
@@
Otherwise, $a=x_{1}a^{\prime}$ for some $a^{\prime}\in\Lambda^{*}\mathbb{R}^{n}$ which is not necessarily in $A$. We show that this cannot happen. Consider the vector space $V$ of elements of $A^{k}$ that have this form. We can choose a vector space $W\subset A^{n-k}$ such that the wedge product defines a nondegenerate bilinear form on $V\otimes W$. In particular, multiplication by $x_{1}$ defines a bijection $W\to x_{1}W$. Now we can similarly build a vector space $V^{\prime}\subset A^{k-1}$ which is Poincaré dual to $x_{1}W$. The map $V^{\prime}\to V$ defined by multiplication by $x_{1}$ is injective, and hence surjective since $V^{\prime}$ and $V$ have the same dimension. Therefore $a$ is contained in the ideal $(x)$ and $\bar{a}=0$, contradicting our assumption.
Finally, since $x_{1}^{2}=0$, multiplication by $x_{1}$ gives a well-defined surjective map from $A/(x_{1})$ to $(x_{1})$. The above argument also shows that this map is also injective, so these two vector spaces have the same dimension. ∎
Corollary 3.5.
Every elliptic or quasiregularly elliptic manifold $M$ with infinite fundamental group has Euler characteristic zero.
This is also proved by a different method in [PG].
Proof.
Choose an element $x\in H^{1}(M)$. Then this follows immediately from the bijection between $A/(x)$ and $(x)$. ∎
@@
Let $M$ be such a manifold. By passing to a finite cover we may assume that $M$ has fundamental group $\mathbb{Z}^{4}$, $\mathbb{Z}^{2}$ or $\mathbb{Z}$. By Corollary 3.5, $M$ has Euler characteristic zero. If $\pi_{1}(M)\cong\mathbb{Z}^{4}$, then by [Hill, Corollary 3.5.1], $M$ is aspherical and therefore homotopy equivalent to $T^{4}$. It is then homeomorphic to $T^{4}$ by [FJ, Theorem 2.16]. Otherwise, by [Hill, Corollary 6.11.1], $M$ is determined by the total Stiefel–Whitney class of the bundle $S^{2}\to M\to T^{2}$ or $S^{3}\to M\to S^{1}$, respectively. Since this class is of finite order, $M$ is finitely covered by either $T^{2}\times S^{2}$ or $S^{1}\times S^{3}$. ∎
This completes the classification of elliptic and quasiregularly elliptic $4$-manifolds up to finite covers. If $\pi_{1}(M)$ is finite, then $M$ is finitely covered by a simply connected manifold which must come from a finite list of 19 manifolds given in [HP, Appendix A].
4. Invariance of ellipticity
Here we restate and prove Theorem 1.5 and Theorem 1.6.
Theorem 1.6.
Let $X=(T^{n-k}\times S^{k})\mathbin{\#}\Sigma$, where $2\leq k\leq n-1$ and $\Sigma$ is a simply connected rational homology sphere. Then $X$ is elliptic.
Proof.
We will show that there is a map $f:T^{n-k}\times S^{k}\to X$ of positive degree. Since $T^{n-k}\times S^{k}$ is elliptic, this completes the proof.
First, note that $\Sigma\setminus\{\text{pt}\}$ is a simply connected space with finite homology groups. It therefore also has finite homotopy groups, in particular there is an $r$ such that the map $S^{n-1}\to\Sigma\setminus\{\text{pt}\}$ winding $r$ times around the puncture is nullhomotopic.
Now consider a map
$\operatorname{id}\times f_{r}:T^{n-k}\times S^{k}\to T^{n-k}\times S^{k}$
where $f_{r}$ is of degree $r$. Without loss of generality, $X$ is built by cutting out a small neighborhood of a regular value of $\operatorname{id}\times f_{r}$ and attaching $\Sigma$. Then we can choose a topological ball $B\subset S^{k}$ which contains the preimages of this neighborhood, and modify the map on $B$ to get a degree $r$ map $T^{n-k}\times S^{k}\to X$. ∎
Theorem 1.5.
Proof.
Suppose that the $n$-manifold $M$ is elliptic, and let $f:\mathbb{R}^{n}\to M$ be a wrapping map. Suppose that $N$ has the same rational homotopy type as $M$. We use Sullivan’s model of rational homotopy theory (see [SulLong, GrMo, FOT]). The manifolds $M$ and $N$ have the same Sullivan minimal algebra $\mathcal{A}$ and minimal models $m_{M}:\mathcal{A}\to\Omega^{*}M$ and $m_{N}:\mathcal{A}\to\Omega^{*}N$. By the shadowing principle ([PCDF, Theorem 4.1] for simply connected targets, [Hansen, Theorem 3.15] for nilpotent targets) the DGA homomorphism $f^{*}m_{M}:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}$ extends to a DGA homomorphism $\Phi:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}\times[0,1]$ such that
  1. (i)
    The forms in the image of $\Phi$ have bounded operator norm, and
  2. (ii)
    $\Phi|_{t=1}=g^{*}m_{N}$ for some Lipschitz map $g:\mathbb{R}^{n}\to N$.
In particular, since
it follows that $g$ is also a wrapping map. ∎